### v1

# PoTable: Programming Standardly on Table-based Reasoning Like a Human Analyst

Abstract
Table-based reasoning has garnered substantial research interest, particularly in its integration with Large Language Model (LLM) which has revolutionized the general reasoning paradigm. Numerous LLM-based studies introduce symbolic tools (e.g., databases, Python) as assistants to extend human-like abilities in structured table understanding and complex arithmetic computations. However, these studies can be improved better in simulating human cognitive behavior when using symbolic tools, as they still suffer from limitations of non-standard logical splits and constrained operation pools. In this study, we propose PoTable as a novel table-based reasoning method that simulates a human tabular analyst, which integrates a Python interpreter as the real-time executor accompanied by an LLM-based operation planner and code generator. Specifically, PoTable follows a human-like logical stage split and extends the operation pool into an open-world space without any constraints. Through planning and executing in each distinct stage, PoTable standardly completes the entire reasoning process and produces superior reasoning results along with highly accurate, steply commented and completely executable programs. Accordingly, the effectiveness and explainability of PoTable are fully demonstrated. Extensive experiments over three evaluation datasets from two public benchmarks on two backbones show the outstanding performance of our approach. In particular, GPT-based PoTable achieves over 4% higher absolute accuracy than runner-ups on all evaluation datasets.
1 Introduction
Tables are widely applied in various scenarios (e.g., healthcare (Ghasemi & Amyot, 2016), finance (Li et al., 2021)), since they can visually present the core information in various types of scientific documents (e.g., articles, reports, websites) (Embley et al., 2006) through a structured format. With the growing development of AI techniques, there has been an increasing demand for automated table processing, attracting significant attention from both academia and industry (Borisov et al., 2022). Recently, the evolution of Large Language Model (LLM) (Zhao et al., 2023) has raised a brand new prompting paradigm for table processing (Lu et al., 2024). This training-free method facilitates complex understanding and reasoning procedures in table question answering (Pasupat & Liang, 2015), table fact verification (Chen et al., 2020) and other downstream tasks (shown in Figure 1(a)).
Throughout the history of humankind, tools have been regarded as the crystallization of human wisdom and a core factor in social productivity development (Washburn, 1960). This consensus has inspired LLM-based techniques to go a step further in simulating more extensive human behavior, i.e., collaborating with symbolic tools to overcome LLMs’ inherent limitations (Qu et al., 2024). In table processing, two unique challenges have been issued in earlier studies (Lu et al., 2024; Dong & Wang, 2024): (1) Tables are structured in two-dimension, leading to unstable memorization of LLMs trained in next-token prediction mode (Sui et al., 2024). (2) Table-based reasoning inevitably involves logical and arithmetic operations, and LLMs may produce misleading results due to their limited calculation abilities. Nevertheless, with a rising trend to utilize databases (Li et al., 2023b), Python (Chen et al., 2022; Gao et al., 2023) and other symbolic tools as assistants, recent approaches effectively reduce table processing errors and misleading computational results by storing the tabular data into internal structure types (e.g., arrays, database tables) and executing syntactic computation commands (e.g., SQL, Python code) (Cheng et al., 2023; Cao et al., 2023; Nahid & Rafiei, 2024).
Despite promising results of the existing studies, there is still room for improvement in simulating human cognitive behavior in table-based reasoning. As shown in Figure 1(b), the human analyst follows a standard tabular data analysis process, considering all possible open-world operations in each stage to reach the local target and finally solve the given tabular task. In general, all the above stages and operations are implemented through a symbolic tool, such as Python with pandas [^1] https://pandas.pydata.org/ package. However, existing studies ignore the standardization of logical split and only consider a few specialized operations for execution. Specifically, some studies attempt to generate the overall program in one go for execution (Cheng et al., 2023), while a few studies iteratively make decisions from pre-defined operation pools without logical boundary (Wang et al., 2024b). On the other hand, partial studies generate executable SQL or Python code of only sub-table and cell content extraction (Ye et al., 2023; Nahid & Rafiei, 2024), while a few studies only specify several commonly used operations for execution (Wang et al., 2024b). In summary, existing executor-based approaches still suffer from non-standard logical splits and constrained operation pools in tool execution. More effective utilization of symbolic tools is required to achieve better performance.
Figure 1: Illustration of (a) two table-based reasoning tasks evaluated in our study, and (b) an example of human cognitive behavior in tabular data analysis, where the human analyst follows a standard data analysis process considering all possible open-world operations to solve the tabular tasks.
Along this line, we propose PoTable (Programming on Tables Standardly) as a novel table-based reasoning method to simulate a human tabular data analyst, which integrates a Python interpreter as the real-time executor accompanied by an LLM-based operation planner and code generator. In PoTable, we split the entire process into several standard table analysis stages, including initialization, row selection, data type cleaning, reasoning and final answering. In each stage, we set a macroscopic goal and expect PoTable to complete the goal through planning and executing. The planning module is in charge of generating chains including any helpful operations under the stage goal. The executing module is responsible for generating executable code based on the current operation and existing generated code, and calling the runtime to execute it at once. Notably, any raised errors lead to the status rollback of the interpreter and the LLM regeneration of the legal corresponding code with returned error information, otherwise it proceeds to the next operation.
It can be clearly seen that PoTable follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution. As a result, PoTable produces superior reasoning results with highly accurate, steply commented and completely executable code, which demonstrates its effectiveness and explainability. Finally, we conduct extensive experiments over three evaluation datasets from two public benchmarks of table-based reasoning tasks on two backbones. PoTable achieves much more outstanding accuracy results than all LLM-based baselines. In particular, GPT-based PoTable achieves over 4% higher absolute accuracy than runner-ups on all evaluation datasets. All experimental results and analyses validate the strong effectiveness of PoTable. In summary, our main contributions can be listed as follows:
  • We propose PoTable, a novel table-based reasoning method that programs standardly to simulate a human analyst. Our method follows a human-like analysis stage split and extends the constrained operation pool into an open-world space for symbolic tool execution.
  • We integrate a Python interpreter as the real-time executor along with an LLM-based operation planner and code generation. PoTable produces high-quality reasoning results with highly accurate, steply commented and completely executable programs, demonstrating its strong effectiveness and explainability.
  • Experimental results over three evaluation datasets from two public benchmarks of table-based reasoning tasks show the outstanding performance of PoTable.
2 PoTable
2.1 Task Formulation
Our study focuses on two table-based reasoning tasks, i.e., table question answering and table fact verification. Each sample can be represented as $(T,Q,A)$, where $T$ denotes the structured table, $Q$ denotes a question to be answered or a statement to be verified. Given $T$ and $Q$, our goal is to find the answer $A$ in the table question answering task, while in the table fact verification task, we have to decide $A=1$ or $A=0$ indicating whether the statement is true or false, respectively.
2.2 Overview
We propose PoTable (Programming on Tables Standardly), a novel table-based reasoning method shown in Figure 2, where we integrate a Python interpreter as the real-time executor along with an LLM-based operation planner and code generator. In PoTable, the entire analysis process is split into five logical stages: initialization, row selection, data type cleaning, reasoning and final answering. In each distinct stage, the planning module is given a macroscopic stage goal to generate operation chains, while the executing module sequentially generates code for the current operation with the existing code base. In most prior studies, symbolic tools are utilized to obtain intermediate results for certain specified operations of decomposition or computation, as LLMs in reality play a dominant role as the global planner and the main processor. In contrast, PoTable offloads the duty of all operation execution to the Python interpreter installing pandas package, while the LLM is responsible for planning reasoning operations and generating code. Consequently, PoTable better leverages LLMs’ advantages in thinking decomposition and code generation, reducing errors from their disadvantages in structured table processing and arithmetic computations.
Figure 2: Illustration of our propose PoTable, a novel table-based reasoning method that simulates a human analyst. PoTable includes five logical stages: initialization, row selection, data type cleaning, reasoning and final answering. Each stage contains a planning module to generate operation chains, and an executing module to generate operation code for real-time execution.
2.3 Human-like Analysis Stage Split
In tabular analysis, humans follow a standard processing pipeline with several stages, such as initialization, pre-processing, comprehensive analysis and post-processing. Standard and clear logical stage splits enhance reasoning effects and enable humans to draw highly accurate and well-supported conclusions. Inspired by human cognitive behavior in tabular analysis, we split the overall analysis procedure into several standard logical stages. These stages will be implemented through Python code with pandas methods as a common choice of human tabular analysts. Specifically, the overall analysis procedure is split into five stages with their own macroscopic goal:
  • Initialization: Store the table data into pandas.DataFrame object.
  • Row Selection: Remove redundant rows that do not represent distinct records.
  • Data Type Cleaning: Transform the data type of table columns into a suitable form.
  • Reasoning: Conduct flexible reasoning operations that are useful to find the final answer.
  • Final Answering: Print out the final answer as the output of the evaluated sample.
In the above stages, the initialization stage is implemented by executing the pre-defined Python code as import pandas as pd and df = pd.DataFrame(data=..., columns=...), and then the LLM and the Python interpreter collaborate to traverse the other stages through meticulous planning and execution sequentially. The detailed procedure is explained in the next subsection. We posit that this stage division is plausible to align with human cognitive behavior in tabular data analysis, which benefits the production of accurate answers and high-quality programs.
2.4 Planning and Executing
To implement the whole table analysis procedure, we adopt a planning module and an executing module to complete the macroscopic goal in each stage. Such deployments leverage the LLM’s advantage in thinking decomposition and code generation, and enjoy the benefits of robust memorization of structured tables and precise computational results of the symbolic tools simultaneously.
Planning. Inspired by Chain-of-Thought (CoT) (Wei et al., 2022) prompting, the LLM decomposes the stage target into operation chains based on the current status of table df, while the output is always formatted as <START>->[OP.]->[OP.]-$\cdots$-><END> for easy operation extraction. We do not restrict the scope of planned operations but only require the operations to be useful in achieving the stage target even the overall tabular task goal. To prompt the LLM, we adopt a few-shot learning strategy (Brown et al., 2020) with three self-made examples for the planning module.
Executing. Given an operation, the LLM generates code based on the current table status df and the existing code base. For the final answering stage, we adopt few-shot prompting with three self-made examples to obtain the code to print out the answer, while in other stages we adopt zero-shot prompting to generate the code. Next, the generated code is sent to the Python interpreter for real-time execution. Most of the time, the execution is successful and then the table status is updated as the next input from df stored in the Python interpreter. Occasionally, the execution fails as the interpreter raises grammar error information or returns illegal output in the final answering stage. In this case, PoTable will roll back the interpreter to the status before the current execution, and urge the LLM to regenerate suitable code based on the abnormal information.
Consequently, the final answer is obtained from the output of the executed code, instead of the direct LLM response of an LLM query. The overall algorithmic procedure is shown in Algorithm 1.
Input: Table $T$, Question or statement $Q$, LLM $M$ and Python Interpreter $R$
Output: Answer $A$ to the question or statement
1
[…]
5       $\operatorname{stageModule}\leftarrow\operatorname{PoTableBlock(PlanPrompt[stage],CodePrompt[stage])}$
6      $\operatorname{operationList}\leftarrow\operatorname{StageModule.plan}(T,Q,M)$
7      for $\operatorname{operaion}$ $\operatorname{in}$ $\operatorname{operationList}$ do
8             $\operatorname{code}\leftarrow\operatorname{stageModule.codeGen}(T,Q,M,\operatorname{operation},\operatorname{codeBase})$
9            $\operatorname{errorCnt}\leftarrow 0$
10            while $\operatorname{catchError}(R.\operatorname{executeCode(code)})$ $\operatorname{as}$ $\operatorname{error}$ $\operatorname{and}$ $\operatorname{errorCnt}<10$ do
11                   $R.\operatorname{resetEnvironment().executeCode(codeBase)}$
12                  $\operatorname{code}\leftarrow\operatorname{stageModule.codeReGen}(T,Q,M,\operatorname{operation,codeBase,error})$
13                  $\operatorname{errorCnt}\leftarrow\operatorname{errorCnt}+1$
14             end while
15            $\operatorname{codeBase}\leftarrow\operatorname{codeBase}+\operatorname{code}$
16            $T\leftarrow R.\operatorname{getCurrentStatus}(T)$
17       end for
18
19 end for
[…]
return $A$
Algorithm 1 PoTable
2.5 Summarization
According to the detailed procedure, we can conclude that PoTable overcomes two explicit limitations in the previous symbolic tools utilization, which enables our work to make great progress in simulating human cognitive behavior in tabular data analysis: (1) Human-like analysis stage split: PoTable programs on tabular data following a standard logical stage split as a human analyst does. As a result, the LLM and the Python interpreter collaborate to complete the planning and executing with little interference in distinct stages. (2) Open-world operation space: PoTable extends the operation pool of the symbolic tools into an open-world space. If the operation benefits the completion of the target goal and the overall process, it will be possibly considered.
As a result, PoTable produces superior reasoning results with high-quality Python programs. The programs are highly accurate, steply commented and completely executable, since they correspond to clear operations and have experienced real-time execution and validation. These advantages and features demonstrate the strong effectiveness and explainability of PoTable.
3 Experiments
3.1 Experimental Setup
Datasets. We conduct experiments on three evaluation sets of two public benchmarks: WikiTQ (Pasupat & Liang, 2015) and TabFact (Chen et al., 2020). WikiTQ is a benchmark for table question answering, which requires answering the question with a short corpus based on the given table. We conduct experiments over the validation (dev.) set with 2,831 questions and the test set with 4,344 questions as previous studies do, and use the official denotation accuracy for evaluation. TabFact is a benchmark for table fact verification, which requires judging whether the given statement is true or false based on the given table. We conduct experiments over the released small test set with 2,024 statements as previous studies do, and use the binary classification accuracy for evaluation.
Backbones. We select two representative language models as the backbones of PoTable and other baseline approaches in our experiments. Specifically, we choose GPT-4o-mini (2024-07-18) [^2] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/ (GPT) as the closed-source small language model, which is competent and cost-efficient to cover a wide range of downstream tasks. In addition, we choose Llama-3.1-70B-Instruct [^3] https://ai.meta.com/blog/meta-llama-3-1/ (LLAMA) as the open-source LLM for evaluation, which shows strong reasoning capabilities among released foundation models. Please refer to Appendix for the detailed parameter settings of the backbone models.
Baselines. We select four competitive LLM-based approaches as baselines for comparison. Binder (Cheng et al., 2023) is a neural-symbolic framework that maps the reasoning task into a specific program and then executes the program binding LLM as a unified API to extend its grammar coverage and tackle the commands that cannot be executed normally. Dater (Ye et al., 2023) first decomposes the table into sub-evidence with column and row selection through LLM queries and then decomposes the question into simpler sub-questions through intermediate SQL generation, followed by a joint reasoning stage with simplified tables and questions. Chain-of-Table (Wang et al., 2024b) pre-defines several common atomic operations for dynamic selection by the LLM, forming an operation chain to process the table with pre-defined code to simplify the table for the final LLM answer querying. TabSQLify (Nahid & Rafiei, 2024) leverages Text-to-SQL to decompose the table into sub-tables and conduct comprehensive reasoning and answer generation through LLM queries. All these selected approaches are competitive as LLM-based baselines on table-based reasoning.
Implementation Details. To implement PoTable, we carefully design prompting templates for planning and executing modules of each stage. In addition, we respectively prepare three few-shot prompting examples for WikiTQ and TabFact, including query-answer pairs for both operation planning and final answer code generation. Please refer to Appendix for the detailed contents.
3.2 Main Results
Comparison Results. We conduct experiments to compare PoTable with other baselines over three evaluation datasets of WikiTQ and TabFact on GPT and LLAMA backbones. The result table is presented in Table 1. From the main results, it is clear that our PoTable significantly outperforms all other baselines over all evaluation datasets from WikiTQ and TabFact on GPT and LLAMA, respectively. In particular, GPT-based PoTable achieves over 4% higher absolute accuracy than runner-ups on all evaluation datasets, which demonstrates the superior effectiveness of our method.
Table 1: Accuracy results (%) of table-based reasoning approaches on WikiTQ (D denotes dev. set, T denotes test set) and TabFact (S denotes small test set) on GPT-4o-mini (GPT) and Llama-3.1-70B-Instruct (LLAMA). The best results are marked in bold and the second-best results are underlined, while the improvements of PoTable over the runner-ups are recorded in teal.
Approach WikiTQ (D) WikiTQ (T) TabFact (S)
GPT LLAMA GPT LLAMA GPT LLAMA
Binder (ICLR’23) 59.20 50.65 58.86 50.51 84.63 78.16
Dater (SIGIR’23) 56.76 42.78 58.33 43.53 80.98 81.57
Chain-of-Table (ICLR’24) 56.64 62.39 55.60 62.22 84.24 85.62
TabSQLify (NAACL’24) 56.87 55.51 57.02 55.78 78.75 70.70
PoTable (Ours)
63.58
(+4.38)
65.10
(+2.71)
64.73
(+5.87)
65.56
(+3.34)
88.93
(+4.30)
87.06
(+1.44)
To be more specific, we make a more comprehensive analysis of results from all approaches and base models. Firstly, Binder is always the runner-up in GPT-based approaches, while its accuracy drop based on LLAMA is 6%-9%. As Binder contains an important step to generate the whole program for the question, it seems that GPT-4o-mini enjoys a higher ability of code generation in one go. In comparison, PoTable has a standard logical split of the whole tabular analysis process and generates code once for a single operation with error checking, reducing possible blurred and uncleared code in the overall programs. Consequently, this may be one reason that our method achieves significant improvement over Binder and others in accuracy. Secondly, in LLAMA-based approaches, Chain-of-Table is the second-best approach although it has a constrained operation pool for dynamic selection, while its accuracy drop based on GPT is around 6% in WikiTQ and 1.44% in TabFact. Its reasoning performance mainly depends on the LLM’s ability to plan and decompose the operations rather than code generating since the codes for all operations are pre-defined, which may indicate that Llama-3.1-70B-Instruct enjoys a stronger ability to plan and reason. In our PoTable, the two LLM abilities are fully stimulated, integrating the advantages of the open-world operation space of symbolic tools for more flexible planning and executing simultaneously. This may be another reason that our method outperforms Chain-of-Table and others in accuracy.
As a result, PoTable overcomes the two limitations in the utilization of symbolic tools by following a human-like logical stage split and executing with an open-world operation space. Compared with the baselines, the way of standard programming in PoTable leads to its best accuracy, strongly validating the effectiveness in table-based reasoning scenarios.
3.3 Performance Analysis Grouped by Task Difficulty and Table Size
To make further performance analysis of PoTable, we recompute the main performance at different levels of task difficulty and table size. Specifically, we label the difficulty of the evaluated questions or statements as “simple” or “complex”. In WikiTQ, a question with a length less than 50 is labeled as “simple”, while a “complex” question is longer. As for TabFact, we use the official difficulty label for all statements. In addition, we group the table content size as “small” (S), “medium” (M) and “large” (L), in situations when the table has 1-49 cells, 50-99 cells and no less than 100 table content cells respectively. The detailed grouped results are reported in Table 2.
The results illustrate that more complex tasks always lead to performance drop as expected, yet the negative correlation between table size and performance is not always obvious. We can draw two preliminary inferences: (1) PoTable may ignore task decomposition as a potential improvement, although it is trivial to see performance drop on difficult tasks. (2) PoTable seems somewhat robust on the table size, yet a deeper study grouped by table tokens may be more persuasive.
Table 2: Accuracy results (%) of PoTable on different groups in task difficulty as simple and complex and different groups in table size as small (S), medium (M) and large (L).
Backbone Dataset Task Difficulty Table Size Original
Simple Complex S M L
GPT WikiTQ (D) 67.77 60.04 63.00 65.57 62.32 63.58
WikiTQ (T) 68.99 61.12 70.20 66.21 61.61 64.73
TabFact (S) 90.65 87.24 90.59 88.44 88.05 88.93
LLAMA WikiTQ (D) 68.00 62.65 68.41 67.52 61.78 65.10
WikiTQ (T) 68.89 62.74 71.27 67.99 61.50 65.56
TabFact (S) 88.96 85.18 85.71 87.84 87.23 87.06
3.4 Ablation Study on Logical Stages
In PoTable, the overall tabular analysis procedure is split into five stages inspired by standard human cognitive behavior. To validate the effect of the logical stage split, we present an ablation study by adopting different stage splits in the compared settings. Specifically, we compare the original GPT-based PoTable with the following four settings: (1) Only Reasoning: we discard all other unnecessary stages except for initialization, reasoning and final answering. In fact, this setting shows no explicit stage split. (2) Removing Row Selection: we give up checking redundant rows before further processing and reasoning stages, which is commonly regarded as an operation of sub-table data extraction. (3) Removing Data Type Cleaning: we give up checking whether the table column data needs type transformation. As all table columns are stored with an initial type of string, discarding this operation may cause more error execution. (4) Adding Column Selection: we add a new column selection stage to select relative columns before further processing and reasoning stages. This stage has been included in most studies as an operation of sub-table data extraction. The overall results of the ablation study on logical stages are shown in Figure 3.
We can see that the original GPT-based PoTable outperforms all ablated settings in the three evaluation datasets. To be more specific, we make a more comprehensive analysis of results from all settings. Firstly, we focus on the “only reasoning” setting (abbreviated as only-rea. below). Compared with the original setting, only-rea. scores nearly 0.6%-1% less in WikiTQ and around 3% less in TabFact. These results indicate that a standard logical split benefits the tabular analysis process. In addition, only-rea. shows few weaknesses among all other settings in WikiTQ but has more accuracy drop in TabFact. From the task perspective, WikiTQ is about a clear task to answer the question directly, while TabFact asks to judge the statement, containing intermediate reasoning processes to judge the potential sub-facts. Therefore, the logical split for TabFact may be more necessary than WikiTQ, and it is also crucial to make the logical stage split as reasonable as possible.
On the other hand, the results of settings “removing row selection” and “removing data type cleaning” demonstrate the importance of these stages as expected. A more interesting fact is the addition of “column selection” results in worse performance, which is widely adopted in previous approaches. We speculate that the selected backbones are competitive enough to handle full table columns, yet eliminating seemingly irrelevant columns may cause a dilemma, i.e., potentially useful columns are accidentally removed and the LLM reasoner cannot find adequate data for processing. Therefore, the operation of column selection is not suitable to be regarded as a distinct stage.
As a completion, we also recompute the performance results grouped by task difficulty and table size. Please refer to Appendix to check the detailed results and analyses.
Figure 3: Accuracy results (%) in the ablation study of the different logic split employed in PoTable with GPT-4o-mini on three evaluation sets of WikiTQ and TabFact, including only reasoning (only Reason), removing row selection (w/o Row Sel.), removing data type cleaning (w/o Dty. Cle.), adding column selection (w/ Col. Sel.) and the original setting (Original). The best results are marked in bold, while the accuracy drops in all settings are recorded in red.
3.5 Case Study
We conduct a case study of PoTable in Figure 4 by presenting an evaluated sample from WikiTQ (T) with its generated Python program and output answer. The tabular task sample is fed into PoTable, experiencing a standard analysis process including five logical stages. From the complete program present, we notice the planned operations (shown as split comments in the stage block) and high-quality generated code of each operation (matching the former comment) for real-time execution, which allows us to review the whole process precisely and discover the true reason why it leads to right or wrong answers. In addition, the generated operations for execution are open-world as long as they are helpful to solve the stage goal, including some helpful commands like inter-mediate variable definition and finding the correct variable as the correct answer to output. Along with the answer, PoTable produces highly accurate, steply commented and completely executable code. These produced outputs indicate the strong effectiveness and explainability of our pipeline.
Figure 4: A case study of an evaluated sample from WikiTQ (T) wit its generated Python program and output answer, which indicates the effectiveness and explainability of PoTable.
4 Related Work
Table Processing with Language Models. Table processing has been a popular research domain over the past decade. Before the era of LLMs, numerous efforts were made to process tables with pre-trained language models. TaPas (Herzig et al., 2020) extends BERT (Devlin et al., 2019) by conducting masked pre-training with joint encoding of questions and flattening tables. TaBERT (Yin et al., 2020) combines content snapshot and vertical attention based on BERT to obtain joint textual and tabular representations for further understanding. TUTA (Wang et al., 2021) enhances transformers (Vaswani et al., 2017) with structure-aware mechanisms to effectively capture spatial, hierarchical and semantic information. TAPEX (Liu et al., 2022) pre-trains BART (Lewis et al., 2020) on a large synthetic SQL dataset to imitate the SQL executor that better understands tabular structure information. With the development of LLMs, the paradigm of table processing has been deeply revolutionized, especially in tabular data encoding and reasoning. In prompting methods, Sui et al. (2024) designs a benchmark to evaluate the structural understanding capabilities of LLMs, followed by a novel self-augmentation for effective structural prompting. Dater (Ye et al., 2023) and DIN-SQL (Pourreza & Rafiei, 2023) adopt task decomposition for better understanding with simplified queries, while Chain-of-Table (Wang et al., 2024b) defines atomic operations for dynamic selection in CoT prompting. Some other approaches explore training or tuning LLMs as generalists. TableLlama (Zhang et al., 2024) develops an open-source tabular LLM by fine-tuning Llama 2-7B (Touvron et al., 2023) with LongLoRA (Chen et al., 2024b), while Table-LLAVA (Zheng et al., 2024) trains a multi-modal tabular LLM that can handle table images as vision inputs.
Table Processing with Symbolic Tools. Symbolic tools have been widely utilized as assistants to produce more accurate and robust mid-results in LLM-based table reasoning scenarios. Most studies adopt databases and Python as affiliated executors to interact with LLMs. Binder (Cheng et al., 2023) and Cao et al. (2023) parse the tasks into integral SQL or Python programs for further execution, incorporating LLM-assistant APIs to handle abstract code blocks for complete execution. TabSQLify (Nahid & Rafiei, 2024) generates SQL queries to extract sub-tables and executes them to get simplified tables for further LLM reasoning. Some works target boosting the code generation ability for tabular reasoning and other scenarios. TroVE (Wang et al., 2024a) asks the code LLMs to curate reusable high-level functions and use them to write solutions for Python execution on the table question answering and other tasks, while Self-Debugging (Chen et al., 2024a) teaches LLMs to debug their predicted SQL or Python programs on Text-to-SQL (Yu et al., 2018) and other tasks. Recently, research in LLM-based table reasoning has been extended into more sophisticated tools environments and more advanced reasoning tasks. SheetCopilot (Li et al., 2023a) and SpreadsheetBench (Ma et al., 2024) address a novel spreadsheet manipulation task, which maneuvers table analysis software like Microsoft Excel [^4] https://www.microsoft.com/zh-cn/microsoft-365/excel to generate step-by-step solutions for simulated execution. MatPlotAgent (Yang et al., 2024) addresses the task of scientific data visualization, which includes a code agent integrating Matplotlib [^5] https://matplotlib.org/ responsible for generating the code to plot figures from input tables. As a result, symbolic tool utilization has become a crucial component in table processing.
5 Conclusion
In this paper, we focused on effective utilization of symbolic tools in LLM-based tabular reasoning. We proposed PoTable, a novel method integrating a Python interpreter as the real-time executor along with an LLM-based operation planner and code generator. PoTable followed a human-like analysis stage split with an open-world operation space for execution. Through standard programming style, PoTable returned answers with highly accurate, steply commented and completely executable code, demonstrating its strong effectiveness and explainability. Extensive experiments under three evaluation datasets of two benchmarks on different backbones presented a dominating performance of PoTable on table-based reasoning. This study targeted standardized tables in the two fundamental reasoning tasks, yet automatic processing of complicated tabular data (e.g., hierarchical tables, multiple tables) and more advanced reasoning tasks (e.g., spreadsheet manipulation, sophisticated machine learning) remained less studied in LLM-based research. In the future, we will explore more effective ways to make our improved method competent on more complicated table-based reasoning scenarios, simulating more advanced human behavior in tabular analysis.
Reproducibility Statement
All baseline approaches have released the official open-source code and prompts. Specifically, we run Binder from https://github.com/xlang-ai/Binder, Dater from https://github.com/AlibabaResearch/DAMO-ConvAI/tree/main/dater, Chain-of-Table from https://github.com/google-research/chain-of-table, and TabSQLify from https://github.com/mahadi-nahid/TabSQLify.